OpenStudy (anonymous):

The lengths of the sides of a right triangle form three consecutive terms in an arithmetic sequence. Show that the triangle is similar to the 3-4-5 right triangle.

OpenStudy (anonymous):

This isn't clear. :/

OpenStudy (anonymous):

its straight from my book

OpenStudy (anonymous):

oh :/

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

@GOODMAN

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

can any of u help me or are u guys just on my question

OpenStudy (anonymous):

hey ask @.Sam.

OpenStudy (anonymous):

Im sorry i dont understand either, as @Ivanka123 said, its unclear.

OpenStudy (anonymous):

k thanks

OpenStudy (mertsj):

Three consecutive terms of arithmetic sequence: a, a+d, a+2d, If the triangles are similar, the corresponding sides have the same ratio: \[\frac{a}{3}=\frac{a+d}{4}=\frac{a+2d}{5}\]

OpenStudy (mertsj):

Can you take it from there?

OpenStudy (anonymous):

I think you might have confused him @Mertsj I feel like this question should contain an image. Like a visual reference to help guide you to the answer.

OpenStudy (mertsj):

That is an excellent idea...could you post one?

OpenStudy (anonymous):

You want me to post one? It seems like the student we're helping is offline. I'll see if I can come up with an image judging by the question.

OpenStudy (mertsj):

Good for you. I guess we care more about the problem than the asker!!

OpenStudy (anonymous):

May this be of assistance to the question?

OpenStudy (mertsj):

Good job. Thanks

OpenStudy (anonymous):

i am sorry i had baseball to go to

OpenStudy (mertsj):

Did you win?

OpenStudy (anonymous):

practice

OpenStudy (mertsj):

What position do you play?

OpenStudy (anonymous):

pitcher

OpenStudy (mertsj):

Are you any good?

OpenStudy (anonymous):

i would say so. can u pls help me with the problem

OpenStudy (anonymous):

Alright, let's see if we can get this problem done.

OpenStudy (mertsj):

I thought I did. Did you read the stuff I posted. It should be almost done.

OpenStudy (mertsj):

Ok. Kitt wants to help you. If you need me you can holler.

OpenStudy (anonymous):

what do i do with the ratios

OpenStudy (anonymous):

@Mertsj please explain the ratios

OpenStudy (mertsj):

You remember that the problem says to show that the triangle is similar to a 3-4-5 right triangle. So we remember that if two triangles have sides with equal ratios that means the triangles are similar.

OpenStudy (mertsj):

So we have two triangles:

OpenStudy (mertsj):